Suppose your firm has decided to use a divisional WACC approach to analyze projects. The firm currently has four divisions, A through D, with average betas for each division of 0.7, 1.0, 1.4, and 1.6, respectively. Assume all current and future projects will be financed with 60 debt and 40 equity, the current cost of equity (based on an average firm beta of 1.0 and a current risk-free rate of 7 percent) is 15 percent and the after-tax yield on the company’s bonds is 10 percent.

What will the WACCs be for each division?
Answer #1

Cost of equity = Rf+ Beta*Risk premium

15%= 7%+1*Rp

Rp = 8%

Division A: Cost of equity = 7%+0.7*8%=12.6%

WACC= 60%*10%+40%*12.6% = 11.04%

Division B: Cost of equity = 7%+1*8%=15%

WACC= 60%*10%+40%*15% = 12%

Division C: Cost of equity = 7%+1.4*8%=18.2%

WACC= 60%*10%+40%*18.2% = 13.28%

Division D: Cost of equity = 7%+1.6*8%=19.8%

WACC= 60%*10%+40%*19.8% = 13.92%